Circular No. 6480

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 960720 L. Piro, E. Costa, M. Feroci, and M. Cinti, Istituto Astrofisica Spaziale,

CNR, Frascati; F. Frontera, D. Dal Fiume, L. Nicastro, M. Orlandini, and G. Pizzichini, Istituto Tecnologie e Studio delle Radiazioni Extraterrestri, CNR, Bologna; J. Heise and R. Jager, Space Research Organization of the Netherlands, Utrecht; F. Fiore, P. Giommi, A. Matteuzzi, and H. Muller, Beppo-SAX Scientific Data Center, Rome; A. Parmar, Space Science Department, ESTEC; B. Sacco, Istituto Fisica Cosmica ed Applicazioni Informatica, Palermo; S. Molendi, Istituto di Fisica Cosmica e Tecnologie Relative, Milano; and R. C. Butler, Agenzia Spaziale Italiana, Rome, report: “BeppoSAX observed the error box of GRB 960720 (IAUC 6467) during Sept. 3.715–4.708 UT for a total of 46 000 s with the MECS telescope and 26 000 s with the LECS telescope, and discovered a source in both telescopes within the WFC error box at α = 17^{h}29^{m}26^{s}.0, δ = +49^{o}04′35′′ (equinox 2000.0), ∼ 2′.9 off the center of the Beppo-SAX WFC error box and ∼ 1′.6 off the ROSAT source (IAUC 6469). We estimate ∼ 1′ for the systematic error in position reconstruction. BeppoSAX and ROSAT positions are consistent within systematic and statistical errors. The LECS (0.1–2 keV) and MECS (1.5–10 keV) countrates are 0.0010 ± 0.0002 and 0.0021 ± 0.0004 count/s, respectively. Taking into account point-spread-function corrections and the instrumental response, we derive (based on preliminary calibrations) the following fluxes: F (0.1–2 keV) = (1.5 ± 0.6) 10^{−13} erg cm^{−2} s^{−1} and F (2–10 keV) = (2.5 ± 0.5) 10^{−13} erg cm^{−2} s^{−1}. The LECS flux is approximately a factor of three greater than the ROSAT flux, assuming a power-law spectrum with an energy index of 1.3–1.5. The LECS/MECS countrate ratio is consistent with no significant absorption in addition to the galactic one (2.6 × 10^{20} cm^{−2}). No other source brighter than ∼ 1 × 10^{−13} erg cm^{−2} s^{−1} is found within the 10′ WFC error box. At least two other sources of flux between 1 and 3 × 10^{−13} erg cm^{−2} s^{−1} lie at ∼ 15′–20′ from the center of the WFC error box. Assuming a value of (log N – log S) of active galactic nuclei from GINGA background fluctuations and recent ASCA data, we estimate the probability of detecting a source with a flux of 2^{−13} erg cm^{−2} s^{−1} within a 5′ radius as ∼ 10 percent.”

COMET C/1996 R1 (HERGENROTHER-SPAHR) Total visual magnitude estimates: Sept. 9.88 UT, 11.8 (R. J. Bouma,

Groningen, The Netherlands, 0.25-m reflector); 14.83, 11.7 (Bouma); 18.20, 11.4 (J. M. Trigo, Castellon, Spain, 0.17-m reflector).

1996 September 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green